# Visual Writing: Writing by Manipulating Visual Representations of Stories

### v1

@@ 4.1. Entities and Actions View
This event-driven representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 1.c). The actions between the entities are represented as directed edges with a label describing the action in one or two words. By default, the nodes are arranged using a force layout to prevent overlaps, but users can rearrange the position of the entities by dragging them.
4.1.1. Editing the Traits of an Entity
Selecting an entity expands its node to reveal traits related to the entity similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”). The intensity of each trait is rated from 1 to 10. This intensity can be changed by moving the slider beside the trait. This results in editing the text of the story to reflect the new intensity of the trait (Figure 2.a).
4.1.2. Adding and Removing an Entity
Adding an entity is done by double-clicking anywhere on the canvas which opens a text input box to indicate the name of the new entity. Once the name is entered, the entity is created at the location of the click. Removing an entity is done by selecting the entity and then pressing “delete” or “backspace” on the keyboard. This results in updating the text of the story so that the entity is effectively removed (Figure 2.b).
4.1.3. Adding and Removing Actions between Entities
Adding an action is done by creating an edge between two nodes. The direction of the edge indicates the source and target of the action. An action can also be created for the same character by connecting a node to itself. This helps represent actions done by an entity with no target (e.g., walking or thinking). An action can be edited at all times by double-clicking its label and changing its description. Similarly, an action can be removed by being selected and then pressing “delete” or “backspace”. As soon as the edge is created, edited, or deleted, the story text is updated to reflect the change (Figure 2.c).
@@ 4.2. Locations and Entities View
The location view is a spatial representation accessed through a tab (Figure 1.d). It shows all the locations in the story as nodes with their name and an emoji representing them. Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position is conditioned by their location in the story. Entities that go to different locations in the story are duplicated to be represented in all of their locations.
4.2.1. Moving an Entity
@@
4.2.2. Adding a Location
A new location can be added by double-clicking anywhere on the canvas (Figure 3.a). This opens a text input field to specify the name of the location. Once the location is created, entities can be moved to it.
@@ 4.3. Timeline of Events View
The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 1.b). These lines are organized one after the other based on the order in which they are presented in the text. The timeline view is always visible. It allows selecting one or multiple events, in which case the other views are updated to fade out the entities and locations not involved in the selected events. Selecting events also forces future modifications to impact only the selected events.
4.3.1. Reordering Events
Events can be dragged to be reorganized temporally. When an event is dragged, it slides horizontally and can be placed between other events (Figure 4.b-c). Once the event is moved, the story text is updated so that the event happens when indicated in the timeline.
4.3.2. Selecting Events
The timeline also allows finding specific scenes and selecting them to allow precise editing. Hovering over events highlights the corresponding sentence in the story (Figure 4.a) and the corresponding entities and locations in the event and location views. Clicking selects the event, and dragging selects multiple events. Once one or multiple events are selected, any visual edits made in the views will target these selected events. For example, selecting an event and then moving an entity in the location view will update the story text so that the entity moved exactly during the selected event.
@@ 4.4. Bi-directional Editor
The rest of the interface is composed of a text editor on the left, a history tree at the bottom to navigate different versions of the story, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 1.a).
4.4.1. Highlighting Visual Elements and Sentences on Hover
Similar to how hovering over the events in the timeline highlights the sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views. This allows for quickly finding different scenes. Similarly, placing the caret in one sentence will act as if the events in that sentence were selected in the timeline. As such, future edits to the visual representations will be restricted to that sentence.
4.4.2. Updating the Visual Representations from the Text
When the text is updated manually, the visual representations might become desynchronized, showing the story as it was before the modification. In this case, the button to refresh the visuals becomes blue to indicate that the visual model may need to be updated. Clicking it results in re-extracting the information from the text and refreshing the views.
4.4.3. Writing the Story from the Visual Representations
By clicking the “refresh from visuals” button, the story is rewritten from scratch using only the visual representations as reference. As such, the story preserves the same events, locations, and entities, but everything else is rewritten. This allows specifying the skeleton of a story and quickly exploring different ways of phrasing the story.
4.4.4. Track changes
By default, all changes made by editing the visual representations are tracked, such that removed text is struck in red while additions are shown in blue. This allows writers to easily locate and review the changes before accepting (or rejecting) them.
4.4.5. History Tree
To make edits easily reversible and facilitate the exploration of alternatives, the system also implements a history tree capable of storing different versions of the story and returning to them at any time. Every time a modification happens, be it by editing the visual representation or editing the text of the story, a snapshot of the story and visual representation are added to the tree. If the writer selects a previous snapshot, then the story is changed to that snapshot. If, after selecting a previous snapshot, a modification is made to the story, then this creates a new branch to the story. Then, writers can navigate between the branches and different versions by clicking on the nodes of the history tree.
@@ 4.5. Implementation
We implemented the visual writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023). Under the hood, modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model.
@@ 4.5.1. Extracting Entities, Locations, and Events
The visual writing approach requires a robust, exhaustive, and fast extraction of the information needed for the visual representations. In our tests, a naive approach consisting of giving a large language model the whole story and asking to extract the necessary information failed at all three of these requirements. Instead, we force the model to output JSON with the proper schema to make the extraction robust. We check that the schema is correct and prompt the model again if necessary. To make the extraction exhaustive, we proceed in multiple steps by first asking the model to extract the entities and locations in the story. We then ask the model to extract events between these entities and at those locations. Finally, and most importantly, we split the whole story into sentences to make the extraction fast. We then prompt the model with the story, but we ask it to extract only the information within the sentence. This allows us to execute the prompt requests in parallel. As such, a text with 30 sentences will send 30 different smaller prompts to the model. This also gives us an indication of the position of the extracted information, as we know that it comes from the sentence being extracted.
@@ 4.5.2. Engineered Prompts to Edit the Story
The system relies on many engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits. We encourage interested readers to review the source code for more details.
5. User Study
We evaluated the visual writing workflow by conducting a user study using our prototype system. Specifically, the study tried to answer the following questions: Are the visual representations understood and helpful? Can people express their editing intents using visual writing? What is the impact of visual writing on exploration and creativity?
Our study is separated into two parts. The first part isolates the use of the three different views in a controlled setting. Because our approach is compatible with other ways to edit text (e.g., no assistance or LLMs) and our main interest is to measure the usefulness of visual writing, we decided against comparing our approach to a baseline. Instead, we collected qualitative insights through semi-structured interviews and explicitly asked participants to compare visual writing with their own workflow. Representations were also evaluated quantitatively using simple usability questionnaires.
The second part of our study was a free-form creative writing task to let participants appropriate the tool and its features.
5.1. Participants
We recruited 8 creative writers from our social network and word of mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male). All had several years of experience with creative writing, including personal writing (N=7), fiction (N=4), scriptwriting (N=2), and poetry (N=2). On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) during their writing process (Mdn=2), how often they find visuals to help during the writing process (Mdn=3), and how frequently they use AI services to help during their writing activities (Mdn=1). Participants reported using a range of types of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2).
5.2. Apparatus
@@
Participants joined the study remotely via the Zoom conference software. They accessed the experiment website using their own computer. The sessions were audio and screen recorded, and interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions). The sessions were about 45 minutes long. In appreciation for their time, participants received a CAD$15 gift card. The study was approved by our institutional ethics board.
5.3. Design and Procedure
Introduction (~2min)
After being greeted, the participant was asked to complete the consent form and demographics questionnaire. They were then told that they would use a system to help them edit stories.
Video Tutorial ((1 + 3 tasks) 30 seconds)
Before each task, the participant watched a short video demonstrating the features they were about to use in the task. The video used a different text and different tasks to make sure the participant could not just reproduce what happened in the tutorial. Before the first task, the participant watched an additional video presenting the concept of the tool and the organization of the interface.
Story Editing Tasks (3 tasks 5min)
Immediately after watching the video tutorial, the participant were prompted to complete the task. We had three task blocks for each of the visual representation views: (1) entities and actions view tasks required removing an entity and then adding a new entity and an action with the main character; (2) location and entities view tasks required moving an entity to an existing location and then creating a new location and move another entity to it; (3) timeline view tasks required the reordering of events. The instructions for each task were phrased as questions and in a way that forced participants to engage with the story. For example, one of the tasks related to the timeline read “What if Jack loses his hat when he flies into the sky?”, forcing participants to understand enough about the story to know where those events occur and move the action to the right location. The tasks were counter-balanced across participants.
Post-Task Questionnaire (3 tasks 2min)
After each task, the participant first indicated how successful they felt in accomplishing the task on a 5-point semantic differential scale. Then, they answered five questions on a 5-point scale related to usability, understandability, and integration into their workflow.
Free-form Exploration (5-15min)
After accomplishing all the tasks and having a good grasp of the tool, the participant engaged in a free-form exploration. We provided the beginning of a story and asked the participant to continue the story with no instructions other than to “explore what could happen next”.
Post-Study Questionnaire (2min)
After completing the free-form part of the study, the participant answered one last questionnaire, which included the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX.
Semi-structured Interview (10min)
At the end, the participant engaged in a semi-structured interview with the experimenter. After giving their immediate impressions of the tool, they were asked about their experience with respect to the main categories of the CSI, like enjoyment, exploration, expressiveness, immersion, and satisfaction given the effort. They were also asked about each representation and their perceived overall advantages and disadvantages of the system.

### v2

@@ 4.1. Entities and Actions View
This representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 3.b). The actions between the entities are represented as directed edges with a label describing the action in one or two words. By default, the nodes are arranged using a force layout to prevent overlaps, but users can rearrange the position of the entities by dragging them.
4.1.1.
Editing the Traits of an Entity
Selecting an entity opens a menu with traits similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”). The intensity of each trait is rated on a scale of 1 to 10. Moving the slider beneath the trait results in editing the text of the story to reflect the new intensity of the trait (Figure 4.a).
4.1.2.
Adding and Removing an Entity
Double-clicking on the canvas opens a text input box, which, once filled, creates an entity with the given name at the location of the mouse pointer. Selecting an entity and then pressing “delete” or “backspace” on the keyboard removes the entity. This results in updating the text of the story so that the entity is effectively removed (Figure 4.b).
4.1.3.
Adding and Removing Actions between Entities
Creating an edge between two nodes opens a text input field to enter a new action. The direction of the edge indicates the source and target of the action. Connecting a node to itself means the action is for the same character. This helps represent actions done by an entity with no target (e.g., walking or thinking). Double-clicking the label of an action edits it. Similarly, selecting an action and pressing “delete” or “backspace” removes it. As soon as the edge is created, edited, or deleted, the story text is updated to reflect the change (Figure 4.c).
4.1.4.
Overlapping Edges and Animations
To declutter the view, overlapping edges are grouped so that only the first action is shown alongside a counter. Users can cycle through the events by clicking the next and previous buttons. Alternatively, hovering over an entity shows animated dots coming and going from the entity, with the name of the action beside them.
@@ 4.2. Locations and Entities View
The location view is a spatial representation accessed through a tab (Figure 3.c). It displays all the locations in the story as nodes with their name and an emoji representing them. Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position depends on their order of appearance in the story. Entities that go to different locations in the story are duplicated to be represented in all their locations.
4.2.1.
Adding a Location
Double-clicking anywhere on the canvas opens a text input field to create a new location (Figure 5.a). Once the location is created, entities can be moved to it.
4.2.2.
Moving an Entity
@@ 4.3. Timeline of Events View
The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 3.a). These lines are organized one after the other based on the order in which they are presented in the text (temporality). The timeline view is always visible. It allows selecting one or multiple events, in which case the other views are updated to fade out the entities and locations not involved in the selected events. Selecting events also forces subsequent modifications to impact only the selected events.
4.3.1.
Selecting Events
The timeline also allows finding specific scenes and selecting them to allow precise editing. Hovering over events highlights the corresponding sentence in the story (Figure 6.a) and the corresponding entities and locations in the event and location views. Clicking selects the event, and dragging selects multiple events. Once events are selected, future visual edits will target these selected events. For example, selecting an event and then moving an entity in the location view updates the story text so that the entity moved exactly during the selected event.
4.3.2.
Reordering Events
Dragging an event slides it horizontally (Figure 6.b-c). Once moved, the story text updates so that the event happens when indicated in the timeline.
@@ 4.4. Bi-directional Editor
The rest of the interface consists of a text editor on the left, a history tree at the bottom, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 3).
4.4.1.
Highlighting Visual Elements and Sentences on Hover
Similar to how hovering over events in the timeline highlights the corresponding sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views. This helps find scenes. Similarly, placing the caret in one sentence will act as if the events in that sentence were selected in the timeline. As such, subsequent edits to the visual representations will be restricted to that sentence.
4.4.2.
Updating the Visual Representations from the Text
When the text is updated manually, the visual representations might become desynchronized, showing the story as it was before the modification. In this case, the refresh button becomes blue to indicate that the visual model might be outdated. Clicking it re-extracts the information from the text and refreshes the views.
4.4.3.
Writing the Story from the Visual Representations
By clicking the “refresh from visuals” button (arrows between the two views in Figure 3), the story is regenerated from scratch using only the visual representations as reference. The regenerated story preserves the same events, locations, and entities, but everything else is rewritten. This allows for specifying the skeleton of a story and quickly exploring different ways of phrasing it.
4.4.4.
Track changes
Changes made by editing the visual representations are tracked, such that removed text is struck while additions are highlighted in green. This allows writers to locate and review the changes before accepting (or rejecting) them.
4.4.5.
History Tree
To make edits easily reversible and facilitate the exploration of alternatives, the system implements a history tree to store different versions of the story and return to them at any time. When a modification happens, be it by editing the visual representation or editing the text, a snapshot of the story and visual representations is added to the tree. When the writer selects a previous snapshot, the story is reverted to that snapshot. If, after selecting a previous snapshot, a modification is made, then a new branch is created. This allows writers to navigate between branches and versions by clicking on the nodes of the history tree.
@@ 4.5. Implementation
We implemented the visual story-writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023). Modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model. A live demo and the source code are accessible online: https://github.com/m-damien/VisualStoryWriting
@@ 4.5.1. Extracting Entities, Locations, and Events
In our tests, the naïve approach of asking an LLM to extract the necessary information results in slow and incomplete extractions. Instead, we first ask the model to extract the entities and locations in the story. We then split the text into sentences and parallelize requests by giving the model the whole story but asking it to extract only the information within each sentence (section A.1). As such, a text with 30 sentences sends 30 smaller prompts in parallel. This leads to faster and more exhaustive extractions. It also helps associate sentences with events. All requests use a structured JSON output.
@@ 4.5.2. Engineered Prompts to Edit the Story
The system relies on engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits. See the appendix and/or the source code for more details.
5. Study 1: Planning Using the Visualizations
In our initial user study, we assessed how story visualizations aid the planning and reviewing phases of writing (Flower and Hayes, 1981). To ensure the study focused solely on visualization, the prototype only supported the
features. We then compared this read-only version to a baseline text-only interface, asking participants to answer high-level planning and reviewing questions. Because no text had to be produced, we recruited participants with various creative writing experiences to reflect the many potential target users of the tool
5.1. Method
5.1.1. Participants
We recruited 12 participants with various writing experiences from our social network and a mailing list of writers (age range: 20 to 37, M=25; 7 self-identified as male, 4 as female, and 1 preferred not to disclose). On a 5-point Likert scale ranging from 1-never to 5-always, they reported how often they write (Mdn=4), how often they use visuals while writing (Mdn=2), and how often they find visuals to help when writing (Mdn=3). For visual aids, some mentioned using timelines (N=5) and diagrams (N=4). They also reported different writing contexts, including academic (N=9), professional (N=7), and creative (N=6) writing.
5.1.2. Apparatus
@@
Participants joined the study remotely via the Zoom conference software and accessed the experiment website using their own computers. The sessions lasted 1 hour and were audio and screen-recorded. Additionally, interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions). Participants received a CAD$20 gift card for their time. Our institutional ethics board approved the study.
5.1.3. Tasks
Tasks simulated what writers ask themselves when reviewing their craft. We extracted questions from best practices on revising fiction (Madden, 1988; Norton, 2011) and then revised them with a professional creative writing instructor. This led to two open-ended questions per story aspect: characters (“Could characters be combined without changing the outcome of the story?” and “Is any character too passive?”); locations (“Are there any locations which could be removed?” and “Are there moments where the spatial logic is broken or unclear?”); time and rhythm (“Is there a large gap between two actions that make the story progress?” and “Is there a scene you think could take place later/earlier in the story?”); and focalization (“If told from another character’s perspective, how would the story change?” and “What does the main character mention that no other character would?”).
5.1.4. Stories
We picked two short stories from the Tell Me A Story dataset (Huot et al., 2025), which consists of human-written stories. Stories were screened to be of similar length (~700 words) and featuring more than two speaking characters and location shifts.
5.1.5. Procedure
After signing the consent form and completing a demographic questionnaire, participants started with one of two conditions: visual story-writing or text-only. For the visual story-writing condition, they watched a one-minute tutorial and then tried the tool. For both conditions, they read the story fully before starting tasks. Participants had to think and answer questions aloud within five minutes, and then explain how they came up with the answer. After each condition, they completed a raw NASA TLX. At the end, they engaged in a semi-structured interview to share impressions, preferences, and strategies. Tasks, stories, and conditions were counterbalanced across participants.
@@ 5.2. Results
Participants were able to complete the tasks without difficulties in both conditions. However, how they accomplished the task varied. Themes are based on their relevance to our research questions, as per an interpretivist approach to reflexive thematic analysis (Braun and Clarke, 2022).
Visualizations helped confirm, analyze, and explore. All participants mentioned using the visualizations to confirm their intuitions. Visualizations gave the “reassurance that my hunch is correct, and I’m not missing anything when answering the questions.” (P6) as they felt “it was mostly a confidence thing where I felt more sure in what I was saying” (P2).
The visuals also helped “have conversations [about] deeper analysis […] and narrative structures” (P4). For example, participants used visuals to “remove this location simply because nobody was there” (P2), “identify who can know about each other” (P5),“figure out if [an event] was playing a part in the story” (P6), and see “how the entities interact with each other” (P12). P8 added it could be particularly useful “if you were editing or doing some kind of work on a piece of text where you’re like […] has everyone been given as much time to speak as I want them to.”
Sometimes, visuals led to discoveries and reflections. For example, P11 mentioned “it kind of dawned on me by looking at the chronology that […] we actually don’t really get [character]’s motivations”. Similarly, P6 used the location view to “reflect on if [location] were to be changed to a different location, would that really have a significant impact on the story? And I realized that’s probably no.” P7 was “paging through the timeline and seeing how the characters jumped around and kind of weren’t even represented […] it actually showed me […] that it’s like, oh, a lot of moments in this story are sort of disconnected from one another”. P3 added “it does make me curious […] how would a story that features less locations look like”.
Visual-driven search vs. skimming. Participants praised the visuals for their help finding passages. With the visuals, “it was way more tidier, similar kind of information were in similar places […] or visually looked similar” (P5) and “if I just click [an entity], it’ll just highlight the parts where this entity is being talked about […] I don’t have to look at other useless portions” (P12). In particular, participants contrasted this to their workflow with text, where “if I just remember, like vaguely, […] then I can probably pattern-match to the visualization faster than trying to figure out like, what is […] the keyword I could search to find the sequence of text.” (P3). In the text condition, participants mentioned how they “tried to hunt for the dialogues […] where there were quotations, and tried to see how many of them were said by [character]” (P4), and were finding locations by “looking for capitalized nouns […] so there’s always the chance that there was something else […] like, oh, the coast!” (P7).
Impact on cognitive load. Results were inconclusive. Wilcoxon signed-rank did not find a significant difference for mental demand (p=.21), physical demand (p=1), temporal demand (p=.49), performance (p=.38), effort (p=.13), nor frustration (p=.16).
Participants’ comments were also mixed. Some mentioned how the visualization reduced their perceived cognitive load because “[with text] I was trying to visualize the things in my head. But with [visuals], it kind of takes away that mental strain” (P12) and “it has all of the characters and locations listed out. If I were to use command F, I would have to put in whatever I can still remember, which oftentimes might not be the complete set”. Others appreciated being able to focus on other aspects because “the visualization […] lets you offload some of the more concrete stuff so that you could explicitly focus on things like character motivations, whereas I found, for the text-based one, I was constantly reskimming in order to refresh this mental schema” (P11). However, others mentioned that the system added to their cognitive load because “trying to learn [the tool] […] almost added this extra layer where I wasn’t engaging with the text” (P7).
Mental model matches and mismatches. The visuals made sense to some participants because “I feel like my brain does this, […] but maybe not this clearly”. In fact, P9, who had the text condition after the visual condition, mentioned how they “laid down a timeline in [their] head, something akin to what we’ve seen in [the visual condition]” to accomplish the tasks.
Yet, sometimes, there was a clear model mismatch. P8 was the most critical, explaining “I’m sure there’s people who just think like this, and this would be a really useful tool for them, because it is an extension of the way they process information. But for me, it’s like really not”. In fact, P8 mentioned often using visuals when writing but explained their process was different: “I will write those scenes on sticky notes and then I can start to arrange them in an order that gives me a plot […], and then I sometimes will use […] picture boards […] to remind me of a visual idea or concept”.
Some other times, the mismatch happened at the semantic level: “calling it an event for me makes it feel like there’s a plot line that’s happening. But it wasn’t really that.” (P2) and “the timeline is very fine grained […] I wasn’t looking at fine-grained scenes […] but just in terms of more large-scale scenes” (P3).
Complementarity of text and visuals Participants viewed the visualization as complementing the text because “you still need the text, the words, to fully appreciate the story” (P1). In fact, when asked how they did the task, participants reported using the views for 74% of tasks, their memory for 46%, and reading for 43% (adds up to more than 100% because two or more ways could be used). One reason is that the visuals pointed to text that then had to be verified. P4 explained “I was using text for the verification, because there were cases when the visualization was capturing something that could be potentially misleading”. Another reason is that the visualization could not capture everything. In particular, the visuals only represented explicit information: “It is really helpful in terms of managing kind of explicit things, like characters, entities, locations […] But it doesn’t quite capture things like motivations” (P11).
6. Study 2: Editing and Free-form Writing
The vision of visual story-writing is to complement the existing writing workflow. As visuals have shown to be helpful for reflecting on the text, we further explore how visual story-writing can assist writers during the “translating” and revision processes of the cognitive model of writing (Flower and Hayes, 1981). To explore this, we recruited experienced creative writers so that they could give us insights on the possible impact the tool would have on their writing workflows. This exploratory study used the fully-featured prototype and comprised two parts: the first part isolates the use of the three different views in a controlled setting, and the second part is a free-form creative writing task. We were interested in the following questions: Are the visual representations understood and helpful? Can people express their editing intents using visual story-writing? What is the impact of visual story-writing on exploration and creativity?
6.1. Method
6.1.1. Participants
We recruited 8 creative writers from our social network and word-of-mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male). None participated in the first study. All had several years of experience with creative writing, including personal writing (N=7), fiction (N=4), scriptwriting (N=2), and poetry (N=2). On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) (Mdn=2), how often they find visuals to help (Mdn=3), and how frequently they use AI services (Mdn=1) during writing. They reported prior use of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2). Participants are labelled W1 to W8 in the rest of the paper.
6.1.2. Apparatus
Same as 5.1.2 except the sessions were about 45 minutes long and participants received a CAD$15 gift card.
6.1.3. Tasks
The first part had three task blocks for each view: (1) entities and actions view tasks required removing an entity then adding a new entity and action involving the main character; (2) location and entities view tasks required moving an entity to an existing location, creating a new location, and move another entity to it; (3) timeline view tasks required reordering events. Instructions were phrased as questions, forcing participants to engage with the story. For example, a timeline task asked “What if Jack loses his hat when he flies into the sky?”. This required participants to understand the story’s events and move the action to the right location. Tasks were counterbalanced across participants. The second part involved a single writing task using a writing prompt.
6.1.4. Procedure
After completing the consent form and demographics questionnaire, participants watched a 30-second video introducing the tool. Before a task block, they watched a short video demonstrating the relevant features. After a block, participants rated their perceived success on a 5-point semantic differential scale and answered five questions related to usability, understandability, and workflow integration. Following all tasks, participants engaged in a free-form exploration to continue a story with no instructions other than to “explore what could happen next.” At the end, participants completed an exit questionnaire (including the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX) and engaged in a semi-structured interview to gather their experience and perceived advantages and disadvantages of the system.
6.2. Results